Point E is between F and G. H is between F and E. FH = 4 in, EG = 9 in, and FH = HE.
FE=_______

HG=_______

FG=______
1
Expert's answer
2020-10-12T16:12:15-0400

FE = FH + HE


But FH=HE


FE = FH + FH = 2FH

FE = 2 × 4 = 8in


HG = HE + EG


Since FH=HE

=> HE=4in


HG = 4 + 9 =13in


FG = FH +HE +EG


FG = 4 + 4 + 9 =17in